# Counterfactual Explanations for Integer Optimization Problems

### v1

@@ Abstract
Counterfactual explanations (CEs) offer a human-understandable way to explain decisions by identifying specific changes to the input parameters of a base or present model that would lead to a desired change in the outcome. For optimization models, CEs have primarily been studied in limited contexts and little research has been done on CEs for general integer optimization problems. In this work, we address this gap. We first show that the general problem of constructing a CE is $\Sigma_{2}^{p}$-complete even for binary integer programs with just a single mutable constraint. Second, we propose solution algorithms for several of the most tractable special cases: (i) mutable objective parameters, (ii) a single mutable constraint, (iii) mutable right-hand-side, and (iv) all input parameters can be modified. We evaluate our approach using classical knapsack problem instances, focusing on cases with mutable constraint parameters. Our results show that our methods are capable of finding optimal CEs for small instances involving up to $40$ items within a few hours.
@@ 1 Introduction
Automated decision-making increasingly shapes critical aspects of modern society, ranging from applications in medical scheduling and financial planning to supply chain optimization and disaster management. Many of these applications are modeled as integer optimization problems (IPs) and are solved by state-of-the-art methods, such as branch and cut. Despite their widespread use, the reasoning behind optimization-based decisions is often unclear to those affected. For instance, optimization algorithms are used for automated planning of surgeries in hospitals [CDB10, BBD+25]. A patient denied an earlier surgery date may ask for an explanation for this decision. In humanitarian logistics, optimal robust depot locations are calculated to quickly dispatch relief items to a region affected by a disaster [SWDHF21]. Here, the properties of an optimal solution significantly depend on the choice of optimization model and uncertainty set. Officials from a certain region may ask for an explanation why their region did not receive a depot while another region did.
@@
One established way to provide explanations for decisions made by automated systems is based on the paradigm of counterfactual explanations (CEs). A CE is a minimal change in model inputs that would lead to a desired change in outcome. For example, in the aforementioned depot location problem, a CE could be: “If the population in your region were 5% higher and the transportation time to region B were 3% smaller, it would be optimal to open a depot in your region.” Counterfactuals have become a cornerstone of explainable artificial intelligence that offers individuals understandable and actionable insights into automated predictions [VBH+24, MRIB22, MKR+24]. Their role in optimization problems—particularly those involving integrality constraints—has only begun to be explored. First works studied the concept of counterfactuals for binary linear optimization problems where only objective parameters are mutable [KB21, KSB21, KB23], drawing on techniques from constraint programming and inverse optimization. Later, Kurtz et al. [KIBd25] identified three types of counterfactuals of interest for optimization problems—weak, strong, and relative CEs—and analyzed them for linear optimization with continuous decision variables. In [LS25] the authors derive a heuristic for calculating weak CEs based on a penalty alternating direction method. Relative CEs have since been investigated for contextual stochastic optimization [RAF25], and CEs have been used to explain efficiency targets in data envelopment analysis [BRAM24] and to detect minimum number of constraints to make a problem infeasible in constraint programming [GOQ24]. Coming back to the case of integer optimization, however, the aforementioned three types of CEs remain largely unexplored, particularly when mutable parameters appear in the constraints.
@@
Constructing CEs is closely related to a number of existing areas of research in optimization. The most obvious connection is to global sensitivity analysis, but rather than quantifying how outcomes vary under systematic parameter changes, CEs ask how parameters must be minimally changed to induce specified outcome changes. This perspective highlights the relationship to inverse optimization—indeed, one may view CE construction as a generalized form of inverse optimization [CMZ25, Wan13]. It also highlights connections to bilevel optimization: the model structures induced by weak CEs are similar to those in optimistic bilevel formulations [DKPVK15, Dem02, KLLS21], whereas strong CEs naturally correspond to pessimistic bilevel formulations [WTKR13]. Certain special cases connect directly to well-studied problem classes. The case in which only the objective function can be modified subsumes the most classical version of inverse optimization, in which one seeks an objective that makes a given solution optimal. The case in which only the right-hand-side can vary relates both to multiobjective optimization via the restricted value function and to the standard form of mixed integer bilevel linear optimization [Dem02]. Finally, the most general case considered here is equivalent to a bilevel optimization model in which the lower-level problem is a mixed integer bilinear optimization problem.
@@
  • We present computational experiments based on knapsack instances from kplib, showing that optimal CEs can be computed in a few hours for instances containing up to $40$ items.
@@ 2 Definitions
In principle, $\mathcal{H}$ can be any (MILP-representable) set. In the remainder of the paper, we choose $\mathcal{H}$ to be a discrete set for two reasons. First, in most applications, parameter values are naturally discrete. Even if we allow the parameters to take on rational values, we can still scale them to be integral without loss of generality. Second, allowing parameter values to be continuous can result in feasible sets that are open, which leads to difficulties in constructing CEs (see [KIBd25] for a discussion on open feasible regions).
We now define the problem of constructing weak and strong CEs. Given a present problem of the form (PP), a favored solution space $\mathcal{D}$, and a mutable parameter space $\mathcal{H}$, a weak CE is any $(c,A,b)\in\mathcal{H}$ such that some optimal solution of the resulting modified instance lies in $\mathcal{D}$, i.e., fulfills the desired properties.
Definition 1.
A weak counterfactual explanation for (PP) is a point $(c,A,b)\in\mathcal{H}$ that satisfies the weak CE condition:
@@
Definition 2.
A strong counterfactual explanation for (PP) is a point $(c,A,b)\in\mathcal{H}$ that satisfies the strong CE condition:
@@
Example 1.
Consider the following instance of (PP):
@@
Figure 1: Counterfactual regions for a ($P$) with right-hand side $b=3$ and three item $i_{1},i_{2},i_{3}$, the latter two of which have a mutable weight. The favored solution space consists of all all solutions that contain $i_{3}$, i.e., $\mathcal{D}_{1}=\{x\in\{0,1\}^{3}:x_{3}=1\}$. Green and orange regions represent strong and weak CEs, respectively. Each region is labeled with the items that are part of an optimal solution and the optimum solution value. The black O marks the parameter values (A^)($\hat${A}) of the present problem.
@@ 3 Complexity
In this section, we show that the problems of deciding whether a counterfactual explanation exists is complete for $\Sigma_{2}^{p}$ in both the weak and strong cases. We use the oracle-based definition of the polynomial hierarchy, as formalized in [AB07], and the polynomial many-to-one notion of reduction introduced by Karp [Kar72]. This means that to prove membership in $\Sigma_{2}^{p}$, we have to show that whenever a weak/strong CE exists, there is a formal certificate of this fact that can be verified in polynomial time, given an oracle for solving problems in $\mathsf{NP}$}. The assumption that we have access to an oracle for solving problems in $\mathsf{NP}$} means, in particular, that we assume that we have the ability to solve the decision version of an ILP in constant time, and hence, the optimization version of ILP in polynomial time.
@@ 3.1 Weak Counterfactual Explanations
When a weak CE exists, equation (1) suggests both a certificate and a method of verifying it that requires us to solve an ILP. It remains to show that there must always exist such a certificate that is polynomial in the size of the input and that this certificate can be verified in polynomial time.
Theorem 1.
@@
If a weak CE exists, this means that there must exist a point $x^{*}\in\mathbb{Z}^{n}$ and a triple $(c^{*},A^{*},b^{*})\in\mathbb{Z}^{n}\times\mathbb{Z}^{m\times n}\times\mathbb{Z}^{n}$ such that $x^{*}\in\operatorname*{arg\,min}_{x\in\mathcal{X}:A^{*}x\geq b^{*}}(c^{*})^{\top}x$ and that $x^{*}\in\mathcal{D}$. For $\tilde{c}:=(c^{*})^{\top}x^{*}$ we must then have that
$(x^{*},c^{*},A^{*},b^{*})\in\mathcal{Q}:=\{(x,c,A,b)\in\mathcal{X}\cap\mathcal{D}\times\mathcal{H}:Ax\geq b,c^{\top}x=\tilde{c}\}$
Since mixed-integer quadratic programming feasibility is in $\mathsf{NP}$} (see Theorem 1 in [PDM17]), there must exist $(\bar{x},\bar{c},\bar{A},\bar{b})\in\mathcal{Q}$ that has size polynomial in the original input data. We claim that $(\bar{x},\bar{c},\bar{A},\bar{b})\in\mathcal{Q}$ is then a certificate that can be checked in polynomial time, given an oracle for problems in $\mathsf{NP}$}.
@@
where all entries of $(\bar{c},\bar{a},\bar{b})$ are positive and integer. Consider a follower, who desires a solution with objective value at least $K$, and a leader who can delete up to $k$ items from the knapsack. Then, the UBIK consists of deciding if there exists a subset of items $I\subseteq[n]$ with $|I|\leq k$ such that for any feasible knapsack solution using only the items from $n]\setminus I$, the objective value is strictly less than a given target $K$. It was shown in [TCCH24] that UBIK is $\Sigma_{2}^{p}$-complete. The UBIK problem is equivalent to an instance of WCE-FEAS where the present problem is
@@
Remark 1.
For the (weak-CE) where only the right hand side $b$ is mutable, it suffices to enumerate all relevant right hand side values of $b$. Assuming that $\mathcal{H}_{b}=\{b\in\mathbb{Z}:\underline{b}\leq b\leq\bar{b}\}$, we then have to solve the problem
@@
This remark leads to a solution algorithm when the present problem is a knapsack problem. If at the same time the favored solution space is given by a fixed number of knapsack constraints, then both problems we have to solve for every value of $b$ can be solved in pseudopolynomial time, which overall leads to a pseudopolynomial algorithm. The exact computational complexity discussion follows.
Theorem 3.
@@
with $q_{i}\in\mathbb{Z}_{+}^{n}$, $p_{i}\in\mathbb{Z}$ for all $i\in[T]$. Furthermore, assume $\mathcal{H}_{b}=\{b\in\mathbb{Z}:\underline{b}\leq b\leq\bar{b}\}$, then (weak-CE) can be solved in time $\mathcal{O}\left(n\bar{b}^{2}\prod_{i=1}^{T}(p_{i}+1)\right)$.
@@
The second problem can be solved by the classical pseudopolynomial algorithm for multi-dimensional knapsack problems in time $\mathcal{O}\left(nb\prod_{i=1}^{T}(p_{i}+1)\right)$ (see [Fré04]) and this dominates the runtime of solving the classical knapsack problem on the left. Since $b\leq\bar{b}$ and we have to solve the latter problems $\bar{b}$ times, the desired result follows. ∎
@@ 3.2 Strong Counterfactual Explanations
$\min_{x\in\mathcal{X}:Ax\geq b}c^{\top}x=\min_{x\in\mathcal{X}\cap\mathcal{D}:Ax\geq b}c^{\top}x\leq\left(\min_{x\in\mathcal{X}:Ax\geq b,x\notin\mathcal{D}}c^{\top}x\right)-1.$ (3)
@@
The first inequality equation that some solution from $\mathcal{D}$ is indeed optimal for the parameters $(c,a,b)$, just as in the weak CE case. The second inequality ensures that all optimal solutions are contained in $\mathcal{D}$. The result follows directly from the definition of strong CE. ∎
@@
$\{(x_{1},x_{2},c,A,b)\in(\mathcal{X}\cap\mathcal{D})\times\mathcal{X}\times\mathcal{H}:Ax_{1}\geq b,Ax_{2}\geq b,c^{\top}x_{1}=\tilde{c}_{1},c^{\top}x_{1}\leq c^{\top}x_{2}-1\}$
Since MIQP feasibility is in $\mathsf{NP}$}, there must exist $(\bar{x}_{1},\bar{x}_{2},\bar{c},\bar{A},\bar{b})\in\mathcal{R}$ that has size polynomial in the original input data (see Theorem 1 in [PDM17]). We claim that $(\bar{x}_{1},\bar{x}_{2},\bar{c},\bar{A},\bar{b})\in\mathcal{R}$ is then a certificate that can be checked in polynomial time, given an oracle for problems in $\mathsf{NP}$}.
The verification requires checking first whether $\bar{x}_{1}\in\mathcal{X}\cap\mathcal{D}$ and $\bar{x}_{2}\in\mathcal{X}$, which can be done in polynomial time under our assumptions. Second, we need to check both that $\bar{x}_{1}\in\operatorname*{arg\,min}_{x\in\mathcal{X}:\bar{A}x\geq\bar{b}}\bar{c}^{\top}x$ and whether $x_{2}^{*}\in\operatorname*{arg\,min}_{x\in\mathcal{X},x\not\in\mathcal{D}:\bar{A}x\geq\bar{b}}\bar{c}^{\top}x$. Both of these checks can also be done in polynomial time using the available $\mathsf{NP}$} oracle.
Theorem 5.
@@
We use exactly the same reduction as in the proof of Theorem 2 with the same inputs, but we simply interpret those inputs as a description of an instance of SCE-FEAS instead of WCE-FEAS.

### v2

@@ Abstract
Counterfactual explanations (CEs) offer a human-understandable way to explain decisions by identifying specific changes to the input parameters of a base or present model that would lead to a desired change in the outcome. For optimization models, CEs have primarily been studied in limited contexts and little research has been done on CEs for general integer optimization problems. In this work, we address this gap. We first show that the general problem of constructing a CE is $\Sigma_{2}^{p}$-complete even for binary integer programs with just a single mutable constraint. Second, we propose solution algorithms for several of the most tractable special cases: (i) mutable objective parameters, (ii) a single mutable constraint, (iii) mutable right-hand-side, and (iv) all input parameters can be modified. We evaluate our approach using classical knapsack problem instances, focusing on cases with mutable constraint parameters. Our results show that our methods are capable of finding optimal CEs for small instances involving up to $40$ items within a few hours. Additionally, we present experiments on the resource constrained shortest path problem.
@@ 1 Introduction
Automated decision-making increasingly shapes critical aspects of modern society, ranging from applications in medical scheduling and financial planning to supply chain optimization and disaster management. Many of these applications are modeled as integer optimization problems (IPs) and are solved by state-of-the-art methods, such as branch and cut. Despite their widespread use, the reasoning behind optimization-based decisions is often unclear to those affected. For instance, optimization algorithms are used for automated planning of surgeries in hospitals [OR_review_2011, ORHC_review_2025]. A patient denied an earlier surgery date may ask for an explanation for this decision. In humanitarian logistics, optimal robust depot locations are calculated to quickly dispatch relief items to a region affected by a disaster [stienen2021optimal]. Here, the properties of an optimal solution significantly depend on the choice of optimization model and uncertainty set. Officials from a certain region may ask for an explanation why their region did not receive a depot while another region did.
@@
One established way to provide explanations for decisions made by automated systems is based on the paradigm of counterfactual explanations (CEs). A CE is a minimal change in model inputs that would lead to a desired change in outcome. For example, in the aforementioned depot location problem, a CE could be: “If the population in your region were 5% higher and the transportation time to region B were 3% smaller, it would be optimal to open a depot in your region.” Counterfactuals have become a cornerstone of explainable artificial intelligence that offers individuals understandable and actionable insights into automated predictions [Verma2024, maragno2022, maragno2024finding]. Their role in optimization problems–particularly those involving integrality constraints–has only begun to be explored. First works studied the concept of counterfactuals for binary linear optimization problems where only objective parameters are mutable [korikov2021counterfactual, korikov2021counterfactualGDPR, korikov2023objective], drawing on techniques from constraint programming and inverse optimization. Later, [kurtz2024] identified three types of counterfactuals of interest for optimization problems—weak, strong, and relative CEs—and analyzed them for linear optimization with continuous decision variables. In [lefebvre2025computing] the authors derive a heuristic for calculating weak CEs based on a penalty alternating direction method. Relative CEs have since been investigated for contextual stochastic optimization [ramirez2025relative], and CEs have been used to explain efficiency targets in data envelopment analysis [bogetoft2024counterfactual] and to detect minimum number of constraints to make a problem infeasible in constraint programming [gupta2024counterfactual]. Coming back to the case of integer optimization, however, the aforementioned three types of CEs remain largely unexplored, particularly when mutable parameters appear in the constraints.
@@
Constructing CEs is closely related to a number of existing areas of research in optimization. The most obvious connection is to global sensitivity analysis, but rather than quantifying how outcomes vary under systematic parameter changes, CEs ask how parameters must be minimally changed to induce specified outcome changes. This perspective highlights the relationship to inverse optimization—indeed, one may view CE construction as a generalized form of inverse optimization [chan2021inverse, Wang2013]. It also highlights connections to bilevel optimization: the model structures induced by weak CEs are similar to those in optimistic bilevel formulations [dempe2015bilevel, dempefoundations, kleinert2021survey], whereas strong CEs naturally correspond to pessimistic bilevel formulations [wiesemann2013pessimistic]. Certain special cases connect directly to well-studied problem classes. The case in which only the objective function can be modified subsumes the most classical version of inverse optimization, in which one seeks an objective that makes a given solution optimal. The case in which only the right-hand-side can vary relates both to multiobjective optimization via the restricted value function and to the standard form of mixed integer bilevel linear optimization [dempefoundations]. Finally, the most general case considered here is equivalent to a bilevel optimization model in which the lower-level problem is a mixed integer bilinear optimization problem.
@@
  • We present computational experiments based on knapsack instances from kplib, showing that optimal CEs can be computed in a few hours for instances containing up to $40$ items. Additionally, we evaluate applications of CEs regarding resource constraint shortest path problems.
@@ 2 Definitions
In principle, $\mathcal{H}$ can be any (MILP-representable) set. In the remainder of the paper, we choose $\mathcal{H}$ to be a discrete set for two reasons. First, in most applications, parameter values are naturally discrete. Even if we allow the parameters to take on rational values, we can still scale them to be integral without loss of generality. Second, allowing parameter values to be continuous can result in feasible sets that are open, which leads to difficulties in constructing CEs; see [kurtz2024] for a discussion on open feasible regions.
We now define the problem of constructing weak and strong CEs. Given a present problem of the form (1), a favored solution space $\mathcal{D}$, and a mutable parameter space $\mathcal{H}$, a weak CE is any $(c,A,b)\in\mathcal{H}$ such that some optimal solution of the resulting modified instance lies in $\mathcal{D}$, i.e., fulfills the desired properties.
Definition 1
A weak counterfactual explanation for (1) is a point $(c,A,b)\in\mathcal{H}$ that satisfies the weak CE condition:
@@
Definition 2
A strong counterfactual explanation for (1) is a point $(c,A,b)\in\mathcal{H}$ that satisfies the strong CE condition:
@@
Example 1
Consider the following instance of (1):
@@
Figure 1: Counterfactual regions for a (1) with right-hand side $b=3$ and three item $i_{1},i_{2},i_{3}$, the latter two of which have a mutable weight. The favored solution space consists of all all solutions that contain $i_{3}$, i.e., $\mathcal{D}_{1}=\{x\in\{0,1\}^{3}:x_{3}=1\}$. Green and orange regions represent strong and weak CEs, respectively. Each region is labeled with the items that are part of an optimal solution and the optimum solution value. The black O marks the parameter values (A^)($\hat${A}) of the present problem.
@@ 3 Complexity
In this section, we show that the problems of deciding whether a counterfactual explanation exists is complete for $\Sigma_{2}^{p}$ in both the weak and strong cases. We use the oracle-based definition of the polynomial hierarchy, as formalized in [AroBar07], and the polynomial many-to-one notion of reduction introduced by [karp1972reducibility]. This means that to prove membership in $\Sigma_{2}^{p}$, we have to show that whenever a weak/strong CE exists, there is a formal certificate of this fact that can be verified in polynomial time, given an oracle for solving problems in $\mathsf{NP}$}. The assumption that we have access to an oracle for solving problems in $\mathsf{NP}$} means, in particular, that we assume that we have the ability to solve the decision version of an ILP in constant time, and hence, the optimization version of ILP in polynomial time.
@@ 3.1 Weak Counterfactual Explanations
When a weak CE exists, Equation (2) suggests both a certificate and a method of verifying it that requires solving an ILP. It remains to show that there must always exist such a certificate that is polynomial in the size of the input and that this certificate can be verified in polynomial time.
Theorem 1
@@
If a weak CE exists, this means that there must exist a point $x^{*}\in\mathbb{Z}^{n}$ and a triple $(c^{*},A^{*},b^{*})\in\mathbb{Z}^{n}\times\mathbb{Z}^{m\times n}\times\mathbb{Z}^{n}$ such that $x^{*}\in\operatorname*{arg\,min}_{x\in\mathcal{X}:A^{*}x\geq b^{*}}(c^{*})^{\top}x$ and that $x^{*}\in\mathcal{D}$. For $\tilde{c}:=(c^{*})^{\top}x^{*}$ we must then have that $(x^{*},c^{*},A^{*},b^{*})\in\mathcal{Q}$ where
$\mathcal{Q}:=\{(x,c,A,b)\in\mathcal{X}\cap\mathcal{D}\times\mathcal{H}:Ax\geq b,c^{\top}x=\tilde{c}\}$
Since mixed-integer quadratic programming feasibility is in $\mathsf{NP}$} (see Theorem 1 in [pia2017mixed]), there must exist $(\bar{x},\bar{c},\bar{A},\bar{b})\in\mathcal{Q}$ that has size polynomial in the original input data. We claim that $(\bar{x},\bar{c},\bar{A},\bar{b})\in\mathcal{Q}$ is then a certificate that can be checked in polynomial time, given an oracle for problems in $\mathsf{NP}$}.
@@
where all entries of $(\bar{c},\bar{a},\bar{b})$ are positive and integer. Consider a follower, who desires a solution with objective value at least $K$, and a leader who can delete up to $k$ items from the knapsack. Then, the UBIK consists of deciding if there exists a subset of items $I\subseteq[n]$ with $|I|\leq k$ such that for any feasible knapsack solution using only the items from $n]\setminus I$, the objective value is strictly less than a given target $K$. It was shown in [tomasaz2024] that UBIK is $\Sigma_{2}^{p}$-complete. The UBIK problem is equivalent to an instance of WCE-FEAS where the present problem is
@@
Remark 1
For the (3) where only the right hand side $b$ is mutable, it suffices to enumerate all relevant right hand side values of $b$. Assuming that $\mathcal{H}_{b}=\{b\in\mathbb{Z}:\underline{b}\leq b\leq\bar{b}\}$, we then have to solve the problem
@@
This remark leads to a solution algorithm when the present problem is a knapsack problem. If at the same time the favored solution space is given by a fixed number of knapsack constraints, then both problems we have to solve for every value of $b$ can be solved in pseudopolynomial time, which leads to a pseudopolynomial algorithm. The exact computational complexity discussion follows.
Theorem 3
@@
with $q_{i}\in\mathbb{Z}_{+}^{n}$, $p_{i}\in\mathbb{Z}$ for all $i\in[T]$. Furthermore, assume $\mathcal{H}_{b}=\{b\in\mathbb{Z}:\underline{b}\leq b\leq\bar{b}\}$, then (3) can be solved in time $\mathcal{O}\left(n\bar{b}^{2}\prod_{i=1}^{T}(p_{i}+1)\right)$.
@@
The second problem can be solved by the classical pseudopolynomial algorithm for multi-dimensional knapsack problems in time $\mathcal{O}\left(nb\prod_{i=1}^{T}(p_{i}+1)\right)$ (see [freville2004multidimensional]) and this dominates the runtime of solving the classical knapsack problem on the left. Since $b\leq\bar{b}$ and we have to solve the latter problems $\bar{b}$ times, the desired result follows. ∎
@@ 3.2 Strong Counterfactual Explanations
$\min_{x\in\mathcal{X}:Ax\geq b}c^{\top}x=\min_{\begin{subarray}{c}x\in\mathcal{X}\cap\mathcal{D}:Ax\geq b\end{subarray}}c^{\top}x\leq\min_{\begin{subarray}{c}x\in\mathcal{X}\setminus D:\\ Ax\geq b\end{subarray}}c^{\top}x-1.$ (6)
@@
The first equation ensures that some solution from $\mathcal{D}$ is indeed optimal for the parameters $(c,a,b)$, just as in the weak CE case. The second inequality ensures that all optimal solutions are contained in $\mathcal{D}$. The result follows directly from the definition of strong CE. ∎
@@
$\{(x_{1},x_{2},c,A,b)\in(\mathcal{X}\cap\mathcal{D})\times\mathcal{X}\times\mathcal{H}:$
$Ax_{1}\geq b,Ax_{2}\geq b,c^{\top}x_{1}=\tilde{c}_{1},$
$c^{\top}x_{1}\leq c^{\top}x_{2}-1\}$
Since MIQP feasibility is in $\mathsf{NP}$}, there must exist $(\bar{x}_{1},\bar{x}_{2},\bar{c},\bar{A},\bar{b})\in\mathcal{R}$ that has size polynomial in the original input data (see Theorem 1 in [pia2017mixed]). We claim that $(\bar{x}_{1},\bar{x}_{2},\bar{c},\bar{A},\bar{b})\in\mathcal{R}$ is then a certificate that can be checked in polynomial time, given an oracle for problems in $\mathsf{NP}$}.
The verification requires checking first whether $\bar{x}_{1}\in\mathcal{X}\cap\mathcal{D}$ and $\bar{x}_{2}\in\mathcal{X}$, which can be done in polynomial time under our assumptions. Second, we need to check both that $\bar{x}_{1}\in\operatorname*{arg\,min}_{x\in\mathcal{X}:\bar{A}x\geq\bar{b}}\bar{c}^{\top}x$ and whether $x_{2}^{*}\in\operatorname*{arg\,min}_{x\in\mathcal{X},x\not\in\mathcal{D}:\bar{A}x\geq\bar{b}}\bar{c}^{\top}x$. Both of these checks can also be done in polynomial time using the available $\mathsf{NP}$} oracle. ∎
Theorem 5
@@
We use exactly the same reduction as in the proof of Theorem 2 with the same inputs, but we simply interpret those inputs as a description of an instance of SCE-FEAS instead of WCE-FEAS. ∎